What is the sketch of y = (-2)^x on a number plane?
My graphics calculators are showing different solutions or no solutions. -.-

Answer by Alan3354(69443)

(-2)^x is not a continuous function.
(-2)^1 = -2
(-2)^2 = +4
(-2)^3 = -8
(-2)^3.5 = a complex number
----------
Graphing calculators use logs internally for exponential calculations, and a log of -2 causes them problems. You get erratic results, if any.